Solve for x:
2x – 3 = 4
2x = 4 + 3
2x = 7
x= 7/2
x = 3.5

Answer by Osran_Shri(18)

2x – 3 = 4
2x = 4 + 3
2x = 7
x = 7/2
x = 3.5